OpenStudy (anonymous):

The answer is no solution right? Solve the system, using substitution. Write the solution as an ordered pair (m, n). 2m - 3n = 6 and m = 3 - 2n (6, 2) (3, 0) (-3, -4) (0, 2) no solution

OpenStudy (anonymous):

I would turn them into lines:\[ 2m - 3n = 6 \implies 2m = 6 + 3n \implies m = 3+1.5n \\ m = 3 - 2n \]These lines do not have the same slope (one is -2, the other is +1.5), so they aren't parallel. So they should eventually hit for some value of \(n\).

OpenStudy (anonymous):

\[ 3-2n = m = 3+1.5n \implies 3-2n = 3+1.5n \]Solving for \(n\) here should tell you the solution, I believe.

OpenStudy (anonymous):

look i did the work i changed m=y and n=x ok 2x-3y=6 and x=3-2y 2(3-2y)-3y=6 6-4y-3y=6 6-1y=6-6 -1y/-1=0/-1 so tha would make it no solution bcs its not possible to get x after that

OpenStudy (anonymous):

Though the substitution method they mentioned would mean doing this: \[ 2m - 3n = 6 \wedge m = 3 - 2n \implies 2(3-2n) - 3n = 6\\ 2(3-2n) - 3n = 6 \]

OpenStudy (anonymous):

so would i get -1/0?

OpenStudy (anonymous):

Okay so -1y/-1=0/-1 means y = 0

OpenStudy (anonymous):

Then plug that into another equation: x=3-2y x=3-2(0) x=3-0 x=3

OpenStudy (anonymous):

ok ok so that means the answer is 3,0 ?

OpenStudy (anonymous):

x = 3, y = 0 (3, 0)

OpenStudy (anonymous):

thank you for your help:)

OpenStudy (anonymous):

No, because you switched m and n

OpenStudy (anonymous):

(m, n), means m is x and n is y.

OpenStudy (anonymous):

i wrpte 3,0 up there didnt you see?

OpenStudy (anonymous):

Oh, ok